OpenStudy (anonymous):

Given the following system of equations, identify the type of system. y = -x + 4 2x + 2y = 8 consistent inconsistent equivalent

OpenStudy (anonymous):

equivalent both represent the sam line

OpenStudy (anonymous):

with 2 unknowns it becomes easy if you rearrange the equations so they are both in y = mx+ c form the second equation: 2y = -2x + 8 dividing by 2: y = -x+ 4 which is equivalent to equation 1